# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 5. Unipotent invariance
We begin with the following general statement.
@@
Let $G\leq\text{GL}(n,\mathbb{R})$ be a closed subgroup. Consider a continuous curve $\varphi:[0,\infty)\to G$ which is unbounded and definable in a polynomially bounded o-minimal structure such that $\det(\varphi(t))$ is constant in $t$. Suppose further that for all $v\in\mathbb{R}^{n}\smallsetminus\{0\}$, $\lim_{t\to\infty}\varphi(t)v\neq 0$. Then there exists a nontrivial one-parameter unipotent group $P\leq G$ such that if $\Gamma\leq G$ is a discrete subgroup, $g\in G$ and $\nu$ is weak-$\ast$ limit of the measures
@@
then $\nu$ is invariant under $P$.
The main idea behind the above statement is the fact that o-minimal curves exhibit “tangency at infinity”. More precisely, there are “change of speed” maps $h(t)$ with $\lim_{t\to\infty}h(t)=\infty$, such that the limit of the differences $\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}=\rho$ exists. Importantly, under our polynomially boundedness and the non-contraction assumption, those matrices $\rho$ will generate a one-parameter unipotent group, and any limiting measure of averaging along $\varphi$ will be invariant under $\rho$.
Such an idea was used in [Shah1994LimitDO] for averaging along polynomial curves. In the polynomial case, it is possible to use Taylor expansion in order to conclude such change of speed maps. For the o-minimal case, Peterzil and Steinhorn in [PS99] showed that the collection of all possible limits of the form $\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}$, where $h$ is a definable function converging to $\infty$ forms a one-dimensional torsion free group. We will refer to this group as the Peterzil-Steinhorn subgroup, which we define below in Definition 5.3 in a simplified manner which suits our needs. Poulios in his thesis [Poulios_thes] studied further the Peterzil-Steinhorn groups, and he proves that the P.S. group for an unbounded curve definable in polynomially bounded structure is either unipotent or $\mathbb{R}$-diagonalizable. Importantly, Poulios establishes a convenient condition for the Peterzil-Steinhorn group to be unipotent.
@@
[Poulios_thes, Chapter 3] Let $G\leq\text{GL}(n,\mathbb{R})$ be a closed subgroup, and consider an unbounded curve $\varphi:[0,\infty)\to G$ definable in a polynomially bounded o-minimal structure. Then, there exists a unique $r\leq 1$ such that the limit
@@
exists and non-zero. Notice that $M_{\varphi}$ is naturally identified as a tangent vector in the Lie algebra $\mathfrak{g}$ of $G$. We have that $M_{\varphi}$ is nilpotent $\iff$ $r<1$, and $M_{\varphi}$ is $\mathbb{R}$-diagonalizable $\iff r=1$. We let $\mathfrak{p}=\text{Span}_{\mathbb{R}}\{M_{\varphi}\}\leq\mathfrak{g}$, and denote as $P\leq G$ the connected subgroup with Lie-algebra $\mathfrak{p}$. Then:
@@
    $r<1$ $\iff$ for each $s\in\mathbb{R}$, the limit
    (5.4) $\rho(s):=\lim_{t\to\infty}\varphi(h_{r,s}(t))\varphi(t)^{-1},$
    exists. In this case $\rho$ defines an isomorphism $\mathbb{R}\to P$. Here $\mathbb{R}$ denotes the additive group of real numbers.
@@
    $r=1\iff$ for each $s>0$, the limit
    (5.5) $\rho(s):=\lim_{t\to\infty}\varphi(h_{1,s}(t))\varphi(t)^{-1},$
    exists. In this case $\rho$ defines an isomorphism $\rho:\mathbb{R}_{>0}\to P$. Here $\mathbb{R}_{>0}$ denotes the multiplicative group of positive real numbers.
@@
We say that $r\leq 1$ is the P.S. order of an unbounded curve $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ definable in a polynomially bounded o-minimal structure, if $r$ satisfies (5.3), and we call the one-parameter subgroup generated by $\exp(M_{\varphi})$ the P.S. group of $\varphi$.
@@
We note that if $\varphi$ as in Proposition 5.2 is a restriction of one-parameter group to $[0,\infty)$, then $r=0$, $h_{r,s}(t)=t+s$, and $\rho(s)=\varphi(s)$ for all $s\in\mathbb{R}$; that is, $\varphi$ is a restriction of a unipotent one-parameter subgroup.
@@
We now show that when the P.S. group is unipotent, any limiting measure of measures of the form (1.1) will be invariant under the P.S. group. This proof is found in [Shah1994LimitDO] and in [Peterzil2018OminimalFO], and we give it here also for completeness. We note the following elementary calculus lemma.
@@
Notice that for all $r<1$, we have $h^{\prime}_{s,r}(t)=1+O(t^{r-1})$.
@@
Let $G\leq\text{GL}(n,\mathbb{R})$ be a closed subgroup and let $\Gamma\leq G$ be a discrete subgroup. Suppose that $\varphi:[0,\infty)\to G$ is a continuous, unbounded curve, definable in a polynomially bounded o-minimal structure such that the P.S. group is unipotent (namely, (5.3) holds for $r<1$). Let $\rho$ as in (5.4). Then
@@
For $f\in C_{c}(G/\Gamma)$ and an $\epsilon>0$, by uniform continuity of $f$, there is an $T_{\epsilon}$ such that
@@
[…]
$+$ $\frac{1}{T}\int_{T_{\epsilon}}^{T}|f(\varphi(h_{r,s}(t)x)-f(\varphi(t)x)|dt+\frac{\epsilon}{2}.$
Now take $T\to\infty$. The first term goes to zero by boundedness of the function $f$, and the second term goes to zero by Lemma 5.5. ∎
Poulios’ observations were used to establish invariance for polynomially bounded o-minimal curves in nilmanifolds by Peterzil and Starchenko, see [Peterzil2018OminimalFO]. If a curve is contained in a unipotent group, as in [Peterzil2018OminimalFO], then the P.S. group is automatically unipotent. Here, in order to verify that the P.S. group is unipotent when the curve is non-contracting we give the following new observation.
@@
We call a curve $\{\varphi(t)\}\subset\text{GL}(n,\mathbb{R})$ essentially diagonal if there exists a decomposition
@@
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be an unbounded continuous curve definable in a polynomially bounded o-minimal structure. Suppose that $\det(\varphi(t))$ is constant for all large enough $t$. Then, the P.S. group of $\varphi$ is unipotent if and only if the curve $\varphi$ is not essentially diagonal.
@@
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be an unbounded continuous curve definable in a polynomially bounded o-minimal structure. Suppose that the determinant $\det(\varphi(t))$ is constant in $t$ and suppose that $\lim_{t\to\infty}\varphi(t)v\neq 0,$ for all $v\in\mathbb{R}^{n}\smallsetminus\{0\}$. Then, $\varphi$ is not essentially diagonal and in particular, the P.S. group of $\varphi$ is unipotent.
@@
where $f_{i,j}\neq 0$, the first non-zero off-diagonal entry. More precisely, $(i,j)$ is the first non-zero entry among the off-diagonal entries according to the following lexicographic order on $\mathbb{N}^{2}$:
@@
Let $B\leq\text{SL}(n,\mathbb{R})$ be the subgroup of upper-triangular matrices. Let $b:[0,\infty)\to B$ be a definable curve. Recall that each definable function $f(t)$ is either zero for all large enough $t$ or $|f(t)|>0$ for all large $t$. Thus, for all large enough $t$, there exists a unique first non-zero off-diagonal entry in $b(t)$, or $b(t)$ is diagonal for all large $t$. We will refer to this entry as the first non-zero off-diagonal entry of the curve $b(t)$.
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a curve definable in an polynomially bounded o-minimal structure. Then there is a definable curve $b:[0,\infty)\to B$, such that $b(t)$ is either diagonal for all large $t$, or the first non-zero off-diagonal entry $f_{i,j}$ satisfies
@@
    $\deg f_{i,j}>\deg f_{j,j}$,
and importantly,
(5.11) $\varphi(t)=\sigma(t)b(t)C,$
where $C\in\text{SL}(n,\mathbb{R})$ is constant, and $\sigma(t)\in\text{SL}(n,\mathbb{R})$ is convergent as $t\to\infty$.
@@
Using the KAN decomposition, we first write $\varphi(t)\in\text{SL}(n,\mathbb{R})$, as $\varphi(t)=k(t)p(t)$, where $k(t)\in\text{SO}(n,\mathbb{R})$ and $p(t)\in B$ an upper-triangular matrix. Since $k(t)$ is obtained by performing the Gram-Schimdt process on the columns of $\varphi$, we conclude that $k(t)$ is definable. As a consequence, $p(t)$ is definable. Since $k(t)$ is bounded and definable, $\lim_{t\to\infty}k(t)$ exists. For all $t$ large, $p(t)$ is either diagonal or $p(t)$ takes the form of (5.9).
@@
      $f_{i,i}-cf_{i,j}$ is eventually zero. If all off-diagonal entries are eventually 00, then we are done. Otherwise, the first eventually non-zero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
@@
    Then, either all the off-diagonal entries in the resulting matrix are eventually 00, and we are done; or the first eventually non-zero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
@@
First, note that it is enough to prove the statement for $\varphi\subset\text{SL}(n,\mathbb{R})$ since the P.S. group is unchanged if we multiply the curve by a constant invertible matrix from the right. In the notations of Lemma 5.10:
@@
Therefore, by Proposition 5.2, the $r$ corresponding to (5.3) for $\varphi(t)$ and $b(t)$ are the same.
Now, $b(t)$ is either eventually diagonal or upper-triangular and satisfies the conditions of the Lemma 5.10 for the first (eventually) non-zero off-diagonal entry.
If $b(t)$ is eventually diagonal, then for all $s>0$, $\lim_{t\to\infty}b(h_{1,s}(t))b^{-1}(t)$ converges to a diagonal matrix. Namely, the P.S. group is diagonal in this case.
Otherwise, let $(i,j)$ be the first non-zero off-diagonal entry in $b(t)$ (for all large enough $t$). We observe that $(i,j)$-th entry in the matrix $b^{\prime}(t)b(t)^{-1}$ is
@@
Recall by Lemma 5.10 that $r_{i,j}-r_{j,j}>0$. Thus the $(i,j)$-th entry in $tb^{\prime}(t)b(t)^{-1}$ is unbounded. So (5.5) is not satisfied. Hence $r\neq 1$. ∎
@@ 5.2. Unipotent invariance in quotients
Consider a non-contracting curve $\varphi:[0,\infty)\to G:=\mathbf{G}(\mathbb{R})$ definable in a polynomially bounded o-minimal structure. By Theorem 4.5, there exists a subsequence of the measures $\mu_{T_{i},\varphi,x}$ as in (1.1) that converges to a probability measure $\mu$ as $i\to\infty$. By Proposition 5.1, $\mu$ is invariant under a nontrivial unipotent one-parameter subgroup. We will let $W$ be the subgroup generated by all unipotent one-parameter subgroups that preserve $\mu$. Then, using Ratner’s theorem describing finite ergodic invariant measures for $W$, and the linearization technique (which is applicable due to the $(C,\alpha)$-good property for the norm of trajectories of $\varphi(t)$ on representations of $\text{SL}(n,\mathbb{R})$), we will prove that $\varphi(t)$ is contained in a certain algebraic subgroup $N$ such that $W\trianglelefteq N$ and that $H_{\varphi}\subseteq N$. Our goal will be to prove that $H_{\varphi}\subset W$. To achieve this goal, we would like to show that if $\varphi(t)$ is not bounded modulo $W$, then the image of $\varphi$ in $N/W$, which is definable, is, in fact, non-contracting, and hence its corresponding P.S. group in $N/W$ is unipotent. Consequently, we will show that $\mu$ is invariant under a unipotent one-parameter subgroup of $N$ that is not contained in $W$. This will contradict our choice of $W$.
@@
As the following example demonstrates, one cannot expect a non-contracting curve modulo a subgroup to have a unipotent P.S. group. Consider $\varphi(t):=\begin{bmatrix}t&t^{2}\\ 0&1/t\end{bmatrix}$ for $t\geq 1$. Then $\varphi$ is non-contracting in $\text{SL}(2,\mathbb{R})$, and its P.S. group is $U:=\{u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}:s\in\mathbb{R}\}$. Now consider the group $P\leq\text{SL}(2,\mathbb{R})$ of upper triangular matrices. Then, $P/U$ is naturally identified with the group $A\leq\text{SL}(2,\mathbb{R})$ of diagonal matrices. In particular, $\varphi(t)$ modulo $H$ is a diagonal curve. In view of our result below, this is explained by the fact that the hull of the curve is not contained in $P$.
We will prove the following.
@@
Let $\psi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be an unbounded, non-contracting curve definable in a polynomially bounded o-minimal structure. Suppose that $\psi\subseteq H_{\psi}$, where $H_{\psi}$ is the hull of $\psi$. Let $N\leq\text{SL}(n,\mathbb{R})$ be a Zariski closed subgroup such that $H_{\psi}\leq N$, and let $H\trianglelefteq N$ be a closed normal subgroup. Let $q:N\to N/H$ be the natural map.
Assume that the image of $\psi$ in $N/H_{u}$ is unbounded. Then, there exists $r<1$ such that
(5.14) $\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1}=p(s),s\in\mathbb{R},$
where $p(s),s\in\mathbb{R}$ is a (non-trivial) one-parameter subgroup of $N/H$. Here, $h_{s,r}(t)$ is given by (5.1). Importantly, there is a one-parameter unipotent subgroup $U\leq N$ such that
$q(U)=\{p(s):s\in\mathbb{R}\}.$

### v2

@@ 5. Unipotent invariance
We say that a measure $\mu$ on $G/\Gamma$ is invariant under $g\in G$ if $g_{*}\mu=\mu$, where $g_{*}$ denotes the pushforward by the left-translation map by $g$ on $G/\Gamma$. We say that a measure is invariant under a subgroup, if it is invariant under every element of the subgroup.
For the rest of the section, $\mathscr{S}_{\mathrm{pbd}}$ is a fixed polynomially bounded o-minimal structure.
@@
Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup. Consider a continuous, unbounded, $\mathscr{S}_{\mathrm{pbd}}$-definable curve $\varphi:[0,\infty)\to G$ such that $\det(\varphi(t))$ is constant in $t$. Suppose further that for all $v\in\mathbb{R}^{n}\setminus\{0\}$, $\lim_{t\to\infty}\varphi(t)v\neq 0$. Then there exists a nontrivial one-parameter unipotent group $P\subseteq G$ such that if $\Gamma\subseteq G$ is a discrete subgroup, $g\in G$ and $\mu$ is weak-$\ast$ limit of the measures
@@
then $\mu$ is invariant under $P$.
The key idea behind the proof is that o-minimal curves exhibit a kind of “tangency at infinity.” Specifically, one can construct definable change-of-speed functions $h(t)$ that diverge to infinity and such that
$\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}=\rho,$
where $\rho$ is a nontrivial element. Under our assumptions, these limiting elements $\rho$ generate a one-parameter unipotent subgroup, and moreover, it holds that $h^{\prime}(t)=1+o(t^{-\nu})$. The latter derivative asymptotic ensures that any limiting measure of $\mu_{T,\varphi,g\Gamma}$ as $T\to\infty$ is invariant under the subgroup generated by $\rho$; see Lemma 5.6 below.
This idea was previously applied by Shah in [Shah1994LimitDO] in the context of averages along polynomial curves. In the polynomial case, one can construct suitable change-of-speed maps using Taylor expansions. In the o-minimal setting, Peterzil and Steinhorn [PS99] showed that the collection of all possible limits of the form $\lim_{t\to\infty}\varphi(h(t))\varphi(t)^{-1}$, where $h$ is a definable function diverging to $\infty$, forms a one-dimensional torsion-free group. We refer to this group as the Peterzil-Steinhorn subgroup, which we define in a simplified form in Definition 5.3 to suit our needs.
Poulios, in his thesis [Poulios_thes], further studied Peterzil-Steinhorn groups and showed that the group associated with an unbounded curve definable in a polynomially bounded structure is either unipotent or $\mathbb{R}$-diagonalizable. Importantly, Poulios establishes a convenient criterion for when the Peterzil-Steinhorn group is unipotent.
@@
[Poulios_thes, Theorem 3.2.4, Theorem 3.3.1, Corollary 3.4.3] Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup, and let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded curve. Then, there exists a unique $r\leq 1$ such that
@@
with $M_{\varphi}\in\mathfrak{g}\neq\{0\}$, where $\mathfrak{g}$ represents the Lie-algebra of $G$.
It holds that $M_{\varphi}$ is nilpotent $\iff$ $r<1$, and $M_{\varphi}$ is $\mathbb{R}$-diagonalizable $\iff r=1$. Moreover:
@@
    If $r<1$, consider $\rho(s):=\exp(sM_{\varphi}),~{}s\in\mathbb{R}$. Then:
    (5.4) $\rho(s):=\lim_{t\to\infty}\varphi(h_{r,s}(t))\varphi(t)^{-1},\ s\in\mathbb{R}.$
@@
    If $r=1$, let $d(s):=\exp(\log(s)M_{\varphi})$, then:
    $d(s):=\lim_{t\to\infty}\varphi(h_{1,s}(t))\varphi(t)^{-1},\ s>0.$
@@
For an unbounded curve $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ definable in a polynomially bounded o-minimal structure, we say that $r\leq 1$ is the P.S. order of $\varphi$, if $r$ satisfies (5.3). We call the one-parameter subgroup $P$ generated by $\exp(M_{\varphi})$ as the P.S. group of $\varphi$.
@@
In the notations of Proposition 5.2, if $\varphi$ is a restriction of a one-parameter group to $[0,\infty)$, then $r=0$, $h_{r,s}(t)=t+s$, and $\rho(s)=\varphi(s)$ for all $s\in\mathbb{R}$; that is, $\varphi$ is a restriction of a unipotent one-parameter subgroup.
@@
We now show that when the P.S. group is unipotent, any limiting measure of measures of the form (1.1) will be invariant under the P.S. group. This proof is found in [Shah1994LimitDO] and in [Peterzil2018OminimalFO], and we give it here also for completeness. The following elementary calculus lemma will be needed.
@@
For the proof below, notice that for $r<1$, it holds that $h^{\prime}_{s,r}(t)=1+O(t^{r-1})$.
@@
Let $G\subseteq\text{GL}(n,\mathbb{R})$ be a closed subgroup and let $\Gamma\subseteq G$ be a discrete subgroup. Suppose that $\varphi:[0,\infty)\to G$ is a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded curve such that its P.S. group is unipotent (namely, (5.3) holds for $r<1$). Let $\rho(s)=\exp(sM_{\varphi})$, as in (5.4). Then
@@
For $f\in C_{c}(G/\Gamma)$ and an $\epsilon>0$, by uniform continuity of $f$, there is an $T_{\epsilon}$ such for all $t\geq T_{\epsilon}$:
@@
$=O\left(\frac{T_{\epsilon}}{T}\right)+O(\epsilon)+\frac{1}{T}\int_{T_{\epsilon}}^{T}\left(f(\varphi(h_{r,s}(t)x)-f(\varphi(t)x)\right)dt.$
The last term converges to zero by Lemma 5.5. ∎
If a curve is contained in a unipotent group, as in [Peterzil2018OminimalFO], then the P.S. group is automatically unipotent. In an overview, to treat our more general setup, we will distinguish between two type of curves to those that are essentially diagonal (see the definition below), and those that are or not. We will show that if a curve is not essentially diagonal, then its P.S. group is unipotent.
@@
We call a curve $\{\varphi(t)\}_{t\geq 0}\subseteq\text{GL}(n,\mathbb{R})$ essentially diagonal if there exists a decomposition
@@
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded, continuous curve. Suppose that $\det(\varphi(t))$ is constant for all large enough $t$. Then, the P.S. group of $\varphi$ is unipotent if and only if the curve $\varphi$ is not essentially diagonal.
@@
Let $\varphi:[0,\infty)\to\text{GL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, unbounded, continuous curve. Suppose that the determinant $\det(\varphi(t))$ is constant in $t$ and suppose that $\lim_{t\to\infty}\varphi(t)v\neq 0,$ for all $v\in\mathbb{R}^{n}\setminus\{0\}$. Then, $\varphi$ is not essentially diagonal and in particular, the P.S. group of $\varphi$ is unipotent.
@@
where $f_{i,j}\neq 0$, the first nonzero off-diagonal entry. More precisely, $(i,j)$ is the first nonzero entry among the off-diagonal entries according to the following lexicographic order on $\mathbb{N}^{2}$:
@@
Let $B\leq\text{SL}(n,\mathbb{R})$ be the subgroup of upper-triangular matrices. Let $b:[0,\infty)\to B$ be a definable curve. Recall that each definable function $f(t)$ is either zero for all large enough $t$ or $|f(t)|>0$ for all large $t$. Thus, for all large enough $t$, there exists a unique first nonzero off-diagonal entry in $b(t)$, or $b(t)$ is diagonal for all large $t$. We will refer to this entry as the first nonzero off-diagonal entry of the curve $b(t)$.
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable curve. Then there exists a definable curve $b:[0,\infty)\to B$ and a constant matrix $C\in\text{SL}(n,\mathbb{R})$ such that
$\varphi(t)=\sigma(t)b(t)C,$
where $\sigma(t)\in\text{SL}(n,\mathbb{R})$ converges as $t\to\infty$. Furthermore, for sufficiently large $t$, the curve $b(t)$ satisfies one of the following:
  • $b(t)$ is diagonal, or
  • the first nonzero off-diagonal entry $f_{i,j}$ of $b(t)$ satisfies:
@@
      $\deg f_{i,j}>\deg f_{j,j}$.
@@
Using the KAN decomposition, we first write $\varphi(t)\in\text{SL}(n,\mathbb{R})$, as $\varphi(t)=k(t)p(t)$, where $k(t)\in\text{SO}(n,\mathbb{R})$ and $p(t)\in B$ an upper-triangular matrix. In view of the Gram-Schimdt process, $k(t)$ and $p(t)$ are definable. Since $k(t)$ is bounded and definable, $\lim_{t\to\infty}k(t)$ exists. For all $t$ large, $p(t)$ is either diagonal or $p(t)$ takes the form of (5.8).
@@
      $f_{i,i}-cf_{i,j}$ is eventually zero. If all off-diagonal entries are eventually 00, then we are done. Otherwise, the first eventually nonzero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
@@
    Then, either all the off-diagonal entries in the resulting matrix are eventually 00, and we are done; or the first eventually nonzero off-diagonal entry in the resulting matrix has a strictly larger index (according to the lexicographic order), and we go back to step 2.
@@
First, note that it is enough to prove the statement for $\varphi\subseteq\text{SL}(n,\mathbb{R})$ since the P.S. group is unchanged if we multiply the curve by a constant invertible matrix from the right. In the notations of Lemma 5.10:
@@
Therefore, by Proposition 5.2, the $r$ corresponding to (5.3) for $\varphi(t)$ and $b(t)$ are the same. Now, $b(t)$ is either eventually diagonal or upper-triangular and satisfies the conditions of the Lemma 5.10 for the first (eventually) nonzero off-diagonal entry. If $b(t)$ is eventually diagonal, then for all $s>0$, $\lim_{t\to\infty}b(h_{1,s}(t))b^{-1}(t)$ converges to a diagonal matrix. Namely, the P.S. group is diagonal in this case.
Suppose that $b(t)$ is not eventually diagonal. Let $(i,j)$ be the first nonzero off-diagonal entry in $b(t)$ (for all large enough $t$). We observe that $(i,j)$-th entry in the matrix $b^{\prime}(t)b(t)^{-1}$ is
@@
Recall by Lemma 5.10 that $r_{i,j}-r_{j,j}>0$. Thus the $(i,j)$-th entry in $tb^{\prime}(t)b(t)^{-1}$ is unbounded. Hence $r\neq 1$. ∎
@@ 5.2. Unipotent invariance in quotients
Consider a non-contracting curve $\varphi:[0,\infty)\to G:=\mathbf{G}(\mathbb{R})$ definable in $\mathscr{S}_{\mathrm{pbd}}$, and consider the measures $\mu_{T,\varphi,x}$; see (1.1). By Theorem 4.5, there exists a subsequence $\mu_{T_{i},\varphi,x}$ which converges to a probability measure $\mu$, as $i\to\infty$. By Proposition 5.1, $\mu$ is invariant under a nontrivial unipotent one-parameter subgroup. Let $W$ be the subgroup generated by all unipotent one-parameter subgroups that preserve $\mu$. We employ Ratner’s measure classification theorem for the $W$-invariant ergodic measures and the linearization technique to prove that $\varphi(t)$ is contained in a certain algebraic subgroup $N$, such that $W\trianglelefteq N$ and such that $H_{\varphi}\subseteq N$, where $H_{\varphi}$ is the hull of $\varphi$. Our goal will be to prove that $H_{\varphi}\subseteq W$. To achieve this goal, we would like to show that if $\varphi(t)$ is not bounded modulo $W$, then the image of $\varphi$ in $N/W$, which is definable, is, in fact, non-contracting, and hence its corresponding P.S. group in $N/W$ is unipotent. The latter outcome implies that $\mu$ is invariant under a unipotent one-parameter subgroup of $N$, which is not contained in $W$. This contradicts our choice of $W$ as the group generated by all unipotents preserving $\mu$. Before proceeding with the relevant result, we have the following remark which emphasizes that the requirement that $H_{\varphi}\subseteq N$ is important.
@@
Consider $\varphi(t):=\begin{bmatrix}t&t^{2}\\ 0&1/t\end{bmatrix}$ for $t\geq 1$. Then $\varphi$ is non-contracting in $\text{SL}(2,\mathbb{R})$, and its P.S. group is $U:=\{u(s):=\begin{bmatrix}1&s\\ 0&1\end{bmatrix}:s\in\mathbb{R}\}$. Now consider the group $B\leq\text{SL}(2,\mathbb{R})$ of upper triangular matrices. Then, $B/U$ is naturally identified with the group $A\leq\text{SL}(2,\mathbb{R})$ of diagonal matrices. In particular, $\varphi(t)$ modulo $U$ is unbounded, but it is a diagonal curve. Due to the next result, this is explained by the fact that $H_{\varphi}$ is not contained in $B$.
@@
Let $\psi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be an unbounded, non-contracting curve definable in $\mathscr{S}_{\mathrm{pbd}}$. Suppose that $\psi([0,\infty))\subseteq H_{\psi}$, where $H_{\psi}$ is the hull of $\psi$. Let $N\subseteq\text{SL}(n,\mathbb{R})$ be a Zariski closed subgroup such that $H_{\psi}\subseteq N$, and let $H\trianglelefteq N$ be a closed normal subgroup.
Assume that the image of $\psi$ in $N/H_{u}$ is unbounded. Let $q:N\to N/H$ be the natural quotient map. Then, there exists $r<1$ such that
(5.12) $\lim_{t\to\infty}q(\psi(h_{s,r}(t)))q(\psi(t))^{-1}=p(s),\,\forall s\in\mathbb{R},$
where $\{p(s):s\in\mathbb{R}\}$ is a nontrivial one-parameter subgroup of $N/H$. Here, $h_{s,r}(t)$ is given by (5.1). Importantly, there exists a one-parameter unipotent subgroup $\{u(s):s\in\mathbb{R}\}\subseteq N$ such that $p(s)=q(u(s))$ for all $s\in\mathbb{R}$.